Question6 Incorrect o.oo points out of 10.00 P Flag question Independent random samples from two regions in the same area gave the following chemical measurements (ppm). Assume the population distributions of the chemical are mound-shaped and symmetric for these two regions Region I: XnI 12 438 1013 1127 737 491 840 306 402 1155 1075 500 340 Region II: x2; n2 16 778 464 563 610 827 894 476 394 824 387 816 767 479 710 389 826 Let μ be the population mean for and μ2 be the population mean for X2. Find a 90% confidence interval for μ1-P2 Select one: a. -119.75 to 248.25 x O b. -253.50 to 125.00 Oc.-248.25 to 119.7 d-114.49 to 242.99 O e.-125.00 to 253.50
Answer #1

Solution:- option d.
Region 1 x1; n1 = 12

x1 = 702, s1 = 326.8074

region 1 x2; n2 = 16

x1 = 637.75 , s2 = 185.6615

df = n1+n2-2 = 12+16-2 = 26

t = 1.706

99% confidence interval for the population difference mean :
(x1-x2) +/- t*sqrt((s1^2/n1)+(s2^2/n2))
  
= (702-637.75)+/- 1.706*sqrt((326.8074^2/12)+(185.6615^2/16))
= -115.12 , 243.62